# Locally integrable cross sections and their intersection covolume

### v2

@@ 6. Intersection spaces
Let $r\in\mathbb{N}$ be arbitrary and fixed, and let $X$ be a Borel $G$-space with a cross section $Y$. Denote by $X^{\otimes r}$ and $Y^{\otimes r}$ the $r$-fold product of $X$ and $Y$, respectively, and consider the diagonal action of $G$ on $X^{r}$, that we denote by
@@ 7. The intersection covolume and a higher order Kac’s lemma
$I_{\mu}\left(Y\right)=\mu^{\left[r,Y\right]}\left(\mathbf{1}_{Y^{\left[r\right]}}\right)$ $=\iint\nolimits_{G\times Y^{\otimes r}}\rho^{Y}\left(g,g^{-1}.\vec{y}\right)\cdot\mathbf{1}_{Y^{\left[r\right]}}\left(g^{-1}.\vec{y}\right)dm_{G}\left(g\right)d\mu_{Y}^{\otimes r}\left(\vec{y}\right)$
@@
$=\mu^{\left[r,Y\right]}\left(\mathbf{1}_{Z^{\left[r\right]}}\right)=I_{\mu}\left(Z\right),$
@@ 7.1. Equivariant Voronoi tessellation scheme
Recall that a set $P\subset G$ is locally finite if its intersection with every compact set in $G$ in finite. By Struble’s theorem [23], every lcsc group $G$ admits a compatible metric, i.e. a right-invariant proper metric $d$ which induces $G$’s topology. The following proposition was observed by Abért and Mellick [1, §3.3].
@@ 7.2. A higher order Kac’s lemma
and since $\prec$ is a Borel order, this $S$ is a Borel set. Then by [13, Theorem (17.25)], the function $Y\to\mathbb{R}_{\geq 0}$, $y\mapsto m_{G}\left(S_{y}\right)=m_{G}\left(\Theta\left(Y_{y},e_{G}\right)\right)$, is measurable, as required.
@@
$=\sum\nolimits_{g\in Y_{\vec{x}}^{\otimes r}}\mathbf{1}_{\Theta\left(Y_{\vec{x}}\,,e_{G}\right)g^{-1}}\left(g^{-1}\right)=\sum\nolimits_{g\in Y_{\vec{x}}^{\otimes r}}\mathbf{1}_{\Theta\left(Y_{\vec{x}}\,,g\right)}\left(e_{G}\right)=1,\quad\vec{x}\in Y^{\left[r\right]},$
where the last equality follows since $\mathscr{T}\left(Y_{\vec{x}}\right)=\{\Theta\left(Y_{\vec{x}},g\right):g\in Y_{\vec{x}}\}$ is a partition of $G$. It follows that
@@ 7.3. Finiteness of the intersection covolume
whose $Y^{\left[r-1\right]}$-periodization (as we did in the proof of Proposition 5.1) is

### v3

@@ 6. Intersection spaces
Let $r\in\mathbb{N}$ be arbitrary and fixed, and let $X$ be a Borel $G$-space with a cross section $Y$. Denote by $X^{\otimes r}$ and $Y^{\otimes r}$ the $r$-fold product of $X$ and $Y$, respectively, and consider the diagonal action of $G$ on $X^{\otimes r}$, that we denote by
@@ 7. The intersection covolume and a higher order Kac’s lemma
$I_{\mu}^{r}\left(Y\right)=\mu^{\left[r,Y\right]}\left(\mathbf{1}_{Y^{\left[r\right]}}\right)$ $=\iint\nolimits_{G\times Y^{\otimes r}}\rho^{Y}\left(g,g^{-1}.\vec{y}\right)\cdot\mathbf{1}_{Y^{\left[r\right]}}\left(g^{-1}.\vec{y}\right)dm_{G}\left(g\right)d\mu_{Y}^{\otimes r}\left(\vec{y}\right)$
@@
$=\mu^{\left[r,Y\right]}\left(\mathbf{1}_{Z^{\left[r\right]}}\right)=I_{\mu}^{r}\left(Z\right),$
@@ 7.1. Equivariant Voronoi tessellation scheme
Recall that a set $P\subset G$ is locally finite if its intersection with every compact set in $G$ in finite. By Struble’s theorem [23], every lcsc group $G$ admits a compatible metric, that is a right-invariant proper metric $d$ which induces $G$’s topology. The following proposition was observed by Abért and Mellick [1, §3.3].
@@ 7.2. A higher order Kac’s lemma
and since $\prec$ is a Borel order, this $S$ is a Borel set. Then by [14, Theorem (17.25)], the function $Y\to\mathbb{R}_{\geq 0}$, $y\mapsto m_{G}\left(S_{y}\right)=m_{G}\left(\Theta\left(Y_{y},e_{G}\right)\right)$, is measurable, as required.
@@
$=\sum\nolimits_{g\in Y_{\vec{x}}^{\otimes r}}\mathbf{1}_{\Theta\big(Y_{\vec{x}}^{\otimes r}\,,e_{G}\big)g^{-1}}\left(g^{-1}\right)=\sum\nolimits_{g\in Y_{\vec{x}}^{\otimes r}}\mathbf{1}_{\Theta\left(Y_{\vec{x}}^{\otimes r}\,,g\right)}\left(e_{G}\right)=1,\quad\vec{x}\in Y^{\left[r\right]},$
where the last equality follows since $\mathscr{T}\left(Y_{\vec{x}}^{\otimes r}\right)=\{\Theta\big(Y_{\vec{x}}^{\otimes r},g\big):g\in Y_{\vec{x}}\}$ is a partition of $G$. It follows that
@@ 7.3. Finiteness of the intersection covolume
whose $Y^{\left[r-1\right]}$-periodization (as in the proof of Proposition 5.1) is